OpenStudy (anonymous):

please help :/ If w=0.50 m/s, diameter of circle is half a meter then projection on diameter has time Period? a) 1 sec b) 5 sec c) pie sec d) 1.5 sec

OpenStudy (anonymous):

@hartnn @nincompoop @kropot72 @Luigi0210 @matricked @Preetha

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

its omega unit is radius /sec *

OpenStudy (sidsiddhartha):

look velocity is given V=.5 m/s and diameter=1/2 so radius=1/4=R now we know that V=Rw w=V/r=4/2=2 again time period T=(2*PI)/w=2*pi/2=pi so option C

OpenStudy (sidsiddhartha):

do u get ut @sarah786

OpenStudy (mrnood):

@sidsiddhartha - I think she changes w to rad/s , not m/s

OpenStudy (anonymous):

yeah answer is C .

OpenStudy (anonymous):

Thank you so much @sidsiddhartha

OpenStudy (sidsiddhartha):

:)